OpenStudy (anonymous):

Integrating e^(x^2) Just wait...I'll post my thing

OpenStudy (anonymous):

\[\int\limits_{}^{}e^x*e^xdx\] \[u = e^x\] \[\frac{du}{dx}=e^x\] \[dx=\frac{du}{e^x}\] \[\int\limits_{}^{}e^x*u*\frac{du}{e^x}=\int\limits_{}^{}udu\] \[=\frac{u^2}{2}+C\] \[=\frac{(e^x)^2}{2}+C\]

OpenStudy (anonymous):

But then when you differentiate it, you don't get e^(x^2) What's going on?

OpenStudy (anonymous):

e^x^2 does not equal (e^x)*(e^x). (e^x)*(e^x)=e^(2x)

OpenStudy (anonymous):

Here is the solution. Your u substitution went wrong. Should have done u=2x.

OpenStudy (anonymous):

Yeah thanks I realised just after I posted. Thanks